If the measure of systolic blood pressure is normally distributed with a mean of 120 and a standard deviation of 10, find the probability that a randomly selected person will have a systolic blood pressure below 130. Assume systolic blood pressure is normally distributed.


1
Expert's answer
2021-04-15T06:56:52-0400

Solution:

Given, "\mu=130,\sigma=10"

"z=\dfrac{X-\mu }{\sigma}"

"P(X<130)=P(z<\dfrac{130-120 }{10})"

"=P(z<1)=0.84134" [Using z-score table]